xx TestFortranCrayPointers

TestFortranCrayPointers

Erik Schnetter <eschnetter@perimeterinstitute.ca>

October 24 2012

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

7 Interfaces

General Implements: testfortrancraypointers

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 gf 10 gf 01 compact 11 0
11                    10                    01 description 11 Grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 ga 10 ga 01 compact 11 0
11                    10                    01 description 11 Grid array
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 10
11                    10  x  01  size 11 10

11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gs 10 gs 01 compact 11 0
11                    10                    01 description 11 Grid scalar
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

8 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestFortranCrayPointers. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxINITIAL x testfortrancraypointersxmain x test fortran cray pointers x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxINITIAL x testfortrancraypointersxcaller x test fortran cray pointers for passing cctkgh x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function